OpenStudy (anonymous):

A greenhouse that specializes in growing geraniums is divided into sections. The number of geranium pots in each section depends on the number of sprinklers in that section. There are x(7x − 2) pots of red geraniums and (x + 3)2 pots of pink geraniums in a section with x sprinklers. Which simplified expression represents the total number of red and pink geranium pots in a section with x sprinklers? 8x2 + 4x + 9 8x2 + 6x + 9 8x2 + 6x + 7 8x4 + 4x2 + 9

OpenStudy (anonymous):

I think this is a matter of simplify, then add, then simplify some more.

OpenStudy (anonymous):

I've tried to solve it but keep getting none of the multi choice options

OpenStudy (anonymous):

You have two expressions, x(7x − 2) and (x + 3)^2. Simplify both of those independently first.

OpenStudy (anonymous):

okay than after that do I add them together?

OpenStudy (anonymous):

Yes, because it's asking for "the total number of red and pink geranium pots "

OpenStudy (anonymous):

i still end up with 8x^2 - 2x +9 ( not one of the options)

OpenStudy (anonymous):

Interesting.. What did you get for (x+3)^2 ?

OpenStudy (anonymous):

x^2 + 9

OpenStudy (anonymous):

Ah.

OpenStudy (anonymous):

That's what I suspected. You're missing the middle term. (x+3)^2 = (x+3)(x+3) = x^2 +3x +3x + 9.

OpenStudy (anonymous):

Exponents don't distribute like multiplication, you have to expand the binomials and multiply by FOIL (First Outer Inner Last).

OpenStudy (anonymous):

oh okay ..... thanks